OpenStudy (anonymous):

Solve the system x + y = 6 and x - y = 2. What is the value of x? A. 6 B. 2 C. 4 D. 0

jimthompson5910 (jim_thompson5910):

x+y=6 x-y=2 Add the equations to get (x+y)+(x-y)=6+2 (x+x)+(y-y) = 8 2x = 8 x = 8/2 x = 4

OpenStudy (anonymous):

The best way to tackle this problem would be to simply add both equations to obtain 2x=8, x=4. Finally, you plug this solution into any of the two equations above to obtain 4+y=6, y=2.

OpenStudy (anonymous):

thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np, glad to be of help

OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? A. (1, 7) B. (5, 0) C. (2, -1) D. (0, 5)

jimthompson5910 (jim_thompson5910):

y = -3x + 5 2x+5 = -3x + 5 2x+3x = 5-5 5x = 0 x = 0/5 x =0 ------------- y = -3x+5 y = -3(0)+5 y = 0+5 y = 5 So the answers are x = 0 and y = 5 giving the answer to be choice D) (0, 5)

OpenStudy (anonymous):

Thank youu!! I have other questions too. I'm sorry to be a bother to you but you're the only one who is really helping me out. You don't just give the answer, you show how to get that answer. And I'm in a rush to get this stuff done because it has to be done by tomorrow. Could you help me with the rest of the questions I have?

jimthompson5910 (jim_thompson5910):

sure, would love to

jimthompson5910 (jim_thompson5910):

don't mind helping out as long as you're learning and taking notes (that's why i do all the steps)

OpenStudy (anonymous):

thanks soo much!! (: Okay.. next question is: Simplify: (2x + 8)(2x - 8). A. 4x^2 - 16 B. 4x^2 - 64 C. 4^x2 - 16x - 64 D. 4x^2 - 8x + 16

OpenStudy (anonymous):

I am taking notes, definitely!!

jimthompson5910 (jim_thompson5910):

(2x + 8)(2x - 8) (2x)(2x) + (2x)(-8) + (8)(2x) + (8)(-8) ... *** Note: use the FOIL rule here 4x^2 - 16x + 16x - 64 4x^2 - 64 So the answer is choice B ***Side note: this is a difference of squares ***

jimthompson5910 (jim_thompson5910):

that's good, hope it's all making (some) sense